# Visual Description Grounding Reduces Hallucinations and Boosts Reasoning in LVLMs

### v2

@@ Abstract
Large Vision-Language Models (LVLMs) often produce responses that misalign with factual information, a phenomenon known as hallucinations. While hallucinations are well-studied, the exact causes behind them remain underexplored. In this paper, we first investigate the root causes of hallucinations in LVLMs. Our findings reveal that existing mitigation techniques primarily reduce hallucinations for visual recognition prompts—those that require simple descriptions of visual elements—but fail for cognitive prompts that demand deliberate reasoning. We identify the core issue as a lack of true visual perception in LVLMs: although they can accurately recognize visual elements, they struggle to fully interpret these elements in the context of the input prompt and effectively link this recognition to their internal knowledge, which is critical for reasoning. To address this gap, we introduce Visual Description Grounded Decoding (VDGD), a simple, robust, and training-free method designed to enhance visual perception and improve reasoning capabilities in LVLMs. VDGD works by first generating a detailed description of the image and appending it as a prefix to the instruction. During response generation, tokens are sampled based on their KL divergence to the description, favoring candidates with lower divergence. Experimental results on multiple visual reasoning benchmarks and LVLMs demonstrate that VDGD consistently outperforms existing baselines 2% - 33%. Finally, we introduce VaLLu, a benchmark designed for comprehensive evaluation of the cognitive capabilities of LVLMs. Code is available here: https://github.com/Sreyan88/VDGD  [^1] Corresponding Author: sreyang@umd.edu. * denotes equal technical contribution.
@@ 1 Introduction
    We conduct extensive controlled experiments to investigate the causes of hallucinations in LVLMs, categorizing them into distinct types. We demonstrate that existing hallucination mitigation techniques are effective only for visual recognition (VR) prompts—specifically, those that involve object recognition and image descriptions—while proving ineffective for cognitive prompts that require reasoning. We further show that these techniques are biased towards reducing language hallucinations in real-world scenes but fail to address hallucinations in non-real-world contexts and other categories.
@@
    Identifying the core issue, we reveal that LVLMs suffer from a critical visual perception gap. Although these models can recognize visual elements and possess the reasoning skills needed for factual response generation, they struggle to fully perceive the visual information in the context of the prompt, resulting in hallucinated outputs for cognitive reasoning tasks.
@@
    We introduce VDGD, a novel and training-free method that enhances LVLM reasoning by grounding response generation in visual descriptions. We evaluate VDGD on X benchmarks with prompts that require deliberate reasoning and how VDGD significantly outperforms existing techniques, improving performance by 2% - 33% by reducing hallucinations.
@@
    Finally, we introduce VaLLu, a meticulously curated benchmark designed to evaluate the cognitive capabilities of LVLMs. Comprising 1,500 carefully selected instances, VaLLu is constructed by refining existing benchmarks and rigorously removing noisy and misleading examples identified through our thorough analysis. VaLLu emphasizes a balanced mix of diverse cognitive tasks, focusing exclusively on questions that require open-ended responses.
@@ 2 Related Work
Large Vision-Language Models In recent years, Large Vision-Language Models (LVLMs) have seen rapid advancements with the release of numerous new models. These developments have been driven by innovations in model architectures (Liu et al., 2023b; Ye et al., 2023; Wang et al., 2023b), training methods (Liu et al., 2023b), alignment techniques (Chen et al., 2024b), and data augmentation techniques (Liu et al., 2023b; Wang et al., 2024b). To assess the capabilities of these models, the research community has introduced a variety of benchmarks specifically designed to evaluate different aspects, including visual recognition and understanding in diverse scenarios (Liu et al., 2023a; Kim et al., 2022), reasoning (Lu et al., 2023; Yue et al., 2024), knowledge-based information retrieval (Hu et al., 2023), and other specialized tasks.
Reducing Hallucinations to Improve Response Quality: Hallucination mitigation broadly aims to enhance the accuracy of model responses. While numerous studies have proposed methods to reduce hallucinations and improve response accuracy (Sicong Leng et al., 2023; Huang et al., 2023b; Yin et al., 2023b; Zhou et al., 2023b) and evaluated their effectiveness (Wang et al., 2023a; Guan et al., 2023), most focus on mitigating object hallucinations in visual recognition tasks. While these efforts are valuable, they fall short of assessing the true reasoning capabilities of LVLMs. Moreover, despite the growing focus on evaluation and mitigation, the underlying causes of hallucinations remain largely unexplored.
@@ 3.2 Do Mitigation Techniques Generalize Beyond Real-World Scenes?
[width=]vllm.drawio5.pdf
@@ 3.3 Are all hallucinations equal?
(3) Style Hallucinations. These are hallucinations caused by style imitation. Open-source VLLMs are often trained on instruction-tuning data synthesized from proprietary closed-source models. Training on such datasets makes the VLLM learn to mimic the characteristics of responses in the IT dataset (Ghosh et al., 2024; Gudibande et al., 2024) (e.g., lengthy answers, a summary line at the end of the response, etc.). However, owing to the lack of sufficient knowledge to mimic its closed-source counterparts, open-source VLLMs might hallucinate facts while responding (Ghosh et al., 2024).
@@ 5.1 Methodology
! Benchmark Baseline LLaVA-v1 LLaVA-1.5 LLaVA-1.6 mPLUG-Owl2 InternLM-X CogVLM CogVLM2 Qwen2-VL MMMU Vanilla-greedy 1.26 1.35 1.42 1.40 2.01 1.66 2.08 2.39 Vanilla-sampling 1.27 1.44 1.40 1.41 2.05 1.64 2.10 2.35 VDD 1.34 1.52 1.44 1.53 2.22 1.68 2.14 2.64 OPERA 1.30 1.43 1.57 1.64 2.25 1.62 2.21 2.44 Woodpecker 1.32 1.44 1.63 1.61 2.12 1.65 2.14 2.56 LRV 1.29 1.49 1.61 1.58 2.08 1.59 2.18 2.48 LURE 1.31 1.47 1.60 1.64 2.27 1.66 2.27 2.71 PAI 1.42 1.39 1.44 1.56 2.23 1.65 2.29 2.64 HALC 1.40 1.54 1.63 1.65 2.15 1.67 2.24 2.69 VDGD (ours) 1.49 (+18%) 1.62 (+20%) 1.75 (+23%) 1.72 (+23%) 2.39 (+19%) 1.71 (+3%) 2.47 (+19%) 2.91 (+22%) MathVista Vanilla-greedy 1.56 1.65 2.00 1.92 2.56 2.15 2.45 2.97 Vanilla-sampling 1.54 1.68 1.91 1.94 2.52 2.19 2.46 2.95 VDD 1.67 1.68 2.07 2.11 2.59 2.53 2.81 3.19 OPERA 1.64 1.83 2.04 2.04 2.62 2.30 2.53 3.22 Woodpecker 1.72 2.10 2.19 2.13 2.43 2.43 2.59 3.13 LRV 1.68 1.68 2.17 1.99 2.61 2.20 2.62 3.30 LURE 1.59 2.03 2.21 2.07 2.37 2.45 2.49 3.13 PAI 1.68 1.85 1.97 2.08 2.63 2.20 2.60 3.39 HALC 1.71 2.05 2.17 1.97 2.54 2.24 2.77 3.38 VDGD (ours) 1.84 (+18%) 2.19 (+33%) 2.44 (+22%) 2.24 (+17%) 2.88 (+13%) 2.62 (+22%) 2.93 (+20%) 3.59 (+21%) MMBench Vanilla-greedy 2.67 2.89 3.11 2.98 3.58 3.19 3.71 3.93 Vanilla-sampling 2.69 2.92 3.12 3.02 3.59 3.22 3.68 3.95 VDD 2.80 3.03 3.09 3.22 3.93 3.28 3.81 4.20 OPERA 2.81 2.94 3.17 3.10 3.95 3.22 3.93 4.05 Woodpecker 2.83 2.96 3.14 3.07 3.84 3.33 3.83 4.19 LRV 2.77 2.99 3.13 3.23 3.84 3.25 3.77 4.05 LURE 2.72 2.97 3.13 3.26 3.72 3.21 3.82 4.16 PAI 2.75 3.00 3.04 3.16 3.94 3.35 3.97 4.05 HALC 2.73 2.93 3.02 3.18 3.81 3.27 4.04 4.08 VDGD (ours) 2.93 (+10%) 3.12 (+8%) 3.26 (+5%) 3.32 (+11%) 4.08 (+14%) 3.42 (+7%) 4.15 (+12%) 4.31 (+10%) MME Vanilla-greedy 3.32 3.54 3.65 3.49 3.75 3.57 3.64 3.86 Vanilla-sampling 3.34 3.53 3.62 3.48 3.77 3.58 3.66 3.85 VDD 3.46 3.61 3.77 3.59 3.96 3.67 3.95 4.21 OPERA 3.42 3.59 3.82 3.54 3.93 3.60 3.88 4.09 Woodpecker 3.37 3.55 3.76 3.63 3.82 3.59 3.92 4.15 LRV 3.43 3.63 3.78 3.52 3.95 3.61 3.96 3.99 LURE 3.42 3.62 3.87 3.67 3.93 3.72 3.77 4.10 PAI 3.38 3.56 3.81 3.55 3.89 3.69 3.82 4.04 HALC 3.47 3.58 3.83 3.71 4.01 3.59 3.80 4.14 VDGD (ours) 3.59 (+8%) 3.70 (+5%) 3.99 (+9%) 3.82 (+9%) 4.12 (+10%) 3.79 (+6%) 4.09 (+12%) 4.34 (+12%) Oven Vanilla-greedy 2.44 2.66 3.13 2.86 3.35 3.35 3.48 3.64 Vanilla-sampling 2.45 2.65 3.10 2.84 3.33 3.37 3.45 3.61 VDD 2.20 2.44 2.96 2.67 3.10 3.21 3.39 3.59 OPERA 2.50 2.49 3.20 2.92 3.44 3.42 3.53 3.68 Woodpecker 2.46 2.46 3.21 2.90 3.45 3.45 3.52 3.70 LRV 2.51 2.52 3.18 2.88 3.47 3.40 3.57 3.74 LURE 2.49 2.51 3.18 2.91 3.42 3.46 3.55 3.72 PAI 2.54 2.60 3.26 2.95 3.48 3.48 3.60 3.78 HALC 2.52 2.57 3.24 2.94 3.50 3.45 3.61 3.80 VDGD (ours) 2.78 (+14%) 2.84 (+7%) 3.49 (+12%) 3.15 (+10%) 3.68 (+10%) 3.59 (+7%) 3.75 (+8%) 3.92 (+7%) LLaVA-Bench Vanilla-greedy 3.01 3.12 3.23 3.13 4.12 3.68 4.15 4.37 Vanilla-sampling 2.98 3.10 3.27 3.17 4.08 3.71 4.17 4.33 VDD 2.93 3.16 3.18 3.15 3.89 3.57 4.13 4.28 OPERA 3.08 3.18 3.30 3.22 4.16 3.75 4.22 4.40 Woodpecker 3.10 3.22 3.36 3.23 4.21 3.73 4.21 4.39 LRV 3.16 3.19 3.30 3.22 4.27 3.72 4.20 4.49 LURE 3.18 3.26 3.35 3.24 4.25 3.64 4.27 4.51 PAI 3.20 3.28 3.38 3.25 4.30 3.74 4.32 4.54 HALC 3.25 3.27 3.37 3.22 4.24 3.76 4.35 4.53 VDGD (ours) 3.36 (+12%) 3.47 (+11%) 3.52 (+9%) 3.38 (+8%) 4.45 (+8%) 3.96 (+8%) 4.50 (+8%) 4.78 (+9%) MMVET Vanilla-greedy 2.39 2.52 2.72 2.87 3.39 3.18 3.47 3.64 Vanilla-sampling 2.34 2.50 2.70 2.84 3.32 3.35 3.42 3.62 VDD 2.29 2.61 2.67 2.92 3.29 3.42 3.39 3.58 OPERA 2.48 2.67 2.77 3.01 3.37 3.39 3.52 3.69 Woodpecker 2.47 2.64 2.82 3.06 3.40 3.44 3.50 3.72 LRV 2.58 2.56 2.79 2.92 3.38 3.42 3.51 3.75 LURE 2.54 2.63 2.81 3.12 3.41 3.46 3.56 3.71 PAI 2.62 2.60 2.83 3.18 3.37 3.49 3.59 3.78 HALC 2.66 2.70 2.88 3.15 3.44 3.47 3.54 3.76 VDGD (ours) 2.85 (+19%) 3.01 (+19%) 3.23 (+19%) 3.45 (+20%) 3.78 (+12%) 3.57 (+12%) 3.74 (+8%) 3.99 (+10%) VaLLu Vanilla-greedy 1.95 2.03 2.63 2.20 2.66 2.82 3.23 3.45 Vanilla-sampling 1.86 2.01 1.64 2.18 2.70 2.83 3.21 3.40 VDD 1.47 1.55 1.80 1.80 2.32 2.63 3.19 3.42 OPERA 2.04 2.05 2.62 2.26 2.65 2.85 3.22 3.47 Woodpecker 2.01 2.05 2.67 2.23 2.60 2.91 3.28 3.52 LRV 1.98 2.10 2.65 2.19 2.59 2.88 3.32 3.51 LURE 2.03 2.03 2.64 2.24 2.64 2.78 3.28 3.48 PAI 2.04 2.23 2.78 2.34 2.76 2.86 3.31 3.54 HALC 2.02 2.22 2.74 2.32 2.72 2.89 3.29 3.52 VDGD (ours) 2.16 (+11%) 2.64 (+30%) 3.16 (+20%) 2.72 (+24%) 3.45 (+30%) 3.01 (+7%) 3.48 (+8%) 3.67 (+6%)

### v3

@@ Abstract
Large Vision-Language Models (LVLMs) often produce responses that misalign with factual information, a phenomenon known as hallucinations. While hallucinations are well-studied, the exact causes behind them remain underexplored. In this paper, we first investigate the root causes of hallucinations in LVLMs. Our findings reveal that existing mitigation techniques primarily reduce hallucinations for visual recognition prompts—those that require simple descriptions of visual elements—but fail for cognitive prompts that demand deliberate reasoning. We identify the core issue as a lack of true visual perception in LVLMs: although they can accurately recognize visual elements, they struggle to fully interpret these elements in the context of the input prompt and effectively link this recognition to their internal knowledge, which is critical for reasoning. To address this gap, we introduce Visual Description Grounded Decoding (VDGD), a simple, robust, and training-free method designed to enhance visual perception and improve reasoning capabilities in LVLMs. VDGD works by first generating a detailed description of the image and appending it as a prefix to the instruction. During response generation, tokens are sampled based on their KL divergence to the description, favoring candidates with lower divergence. Experimental results on multiple visual reasoning benchmarks and LVLMs demonstrate that VDGD consistently outperforms existing baselines 2% - 33%. Finally, we introduce VaLLu, a benchmark designed for comprehensive evaluation of the cognitive capabilities of LVLMs.
@@ 1 Introduction
    Through extensive experiments, we categorize hallucinations in LVLMs and show that existing mitigation techniques work well for visual recognition prompts but fail for reasoning-intensive cognitive prompts.
@@
    Inspired by this, we identify a critical visual perception gap in LVLMs: while they can recognize visual elements and reason, they struggle to contextualize visual information with the prompt, leading to hallucinations in cognitive reasoning tasks.
@@
    We introduce VDGD, a novel and training-free method that enhances LVLM reasoning by grounding response generation in visual descriptions. We evaluate VDGD on 8 benchmarks with prompts that require deliberate reasoning and how VDGD significantly outperforms existing techniques, improving performance by 2% - 33% by reducing hallucinations.
@@
    Finally, we introduce VaLLu, a meticulously curated benchmark designed to evaluate the cognitive capabilities of LVLMs.
@@ 2 Related Work
Large Vision-Language Models In recent years, Large Vision-Language Models (LVLMs) have seen rapid advancements with the release of numerous new models. These developments have been driven by innovations in model architectures (Liu et al., 2023b; Ye et al., 2023; Wang et al., 2023b), training methods (Liu et al., 2023b), alignment techniques (Chen et al., 2024b), and data augmentation techniques (Liu et al., 2023b; Wang et al., 2024b). To assess the capabilities of these models, the research community has introduced a variety of benchmarks specifically designed to evaluate different aspects, including visual recognition and understanding in diverse scenarios (Liu et al., 2023a; Kim et al., 2022), reasoning (Lu et al., 2023; Yue et al., 2024; Bai et al., 2024; Liu et al., 2024b), knowledge-based information retrieval (Hu et al., 2023), and other specialized tasks.
Reducing Hallucinations to Improve Response Quality: Hallucination mitigation broadly aims to enhance the accuracy of model responses. While numerous studies have proposed methods to reduce hallucinations and improve response accuracy (Sicong Leng et al., 2023; Huang et al., 2023b; Yin et al., 2023b; Zhou et al., 2023b) and evaluated their effectiveness (Wang et al., 2023a; Guan et al., 2023), most focus on mitigating object hallucinations in visual recognition tasks. While these efforts are valuable, they fall short of assessing the true reasoning capabilities of LVLMs. Moreover, despite the growing focus on evaluation and mitigation, the underlying causes of hallucinations remain largely unexplored. Recent methods like Compositional Chain-of-Thought Prompting (CCoT)(Mitra et al., 2024), Visual Table(Zhong et al., 2024), and Visual Evidence Prompting (Li et al., 2024) introduce structured intermediate representations (e.g., scene graphs or hierarchical descriptions) to improve reasoning. While these methods show strong performance on real-world scenes, they rely heavily on explicit object-centric representations and often struggle with abstract or non-real-world inputs (e.g., charts or mathematical graphs). Additionally, methods like Visual Table are not training-free and require substantial computational resources. In contrast, VDGD is lightweight, training-free, and generalizes effectively to diverse inputs, including abstract datasets like MMMU and MathVista.
@@ 3.2 Do Mitigation Techniques Generalize Beyond Real-World Scenes?
[width=]vllm.drawio5-compressed.pdf
@@ 3.3 Are all hallucinations equal?
(3) 1.3 Hallucinations. These are hallucinations caused by style imitation. Open-source VLLMs are often trained on instruction-tuning data synthesized from proprietary closed-source models. Training on such datasets makes the VLLM learn to mimic the characteristics of responses in the IT dataset (Ghosh et al., 2024; Gudibande et al., 2024) (e.g., lengthy answers, a summary line at the end of the response, etc.). However, owing to the lack of sufficient knowledge to mimic its closed-source counterparts, open-source VLLMs might hallucinate facts while responding (Ghosh et al., 2024).
@@ 5.1 Methodology
! Benchmark Baseline LLaVA-v1 LLaVA-1.5 LLaVA-1.6 mPLUG-Owl2 InternLM-X CogVLM CogVLM2 Qwen2-VL MMMU Vanilla-greedy 1.26 1.35 1.42 1.40 2.01 1.66 2.08 2.39 Vanilla-sampling 1.27 1.44 1.40 1.41 2.05 1.64 2.10 2.35 VCD 1.34 1.52 1.44 1.53 2.22 1.68 2.14 2.64 OPERA 1.30 1.43 1.57 1.64 2.25 1.62 2.21 2.44 Woodpecker 1.32 1.44 1.63 1.61 2.12 1.65 2.14 2.56 LRV 1.29 1.49 1.61 1.58 2.08 1.59 2.18 2.48 LURE 1.31 1.47 1.60 1.64 2.27 1.66 2.27 2.71 PAI 1.42 1.39 1.44 1.56 2.23 1.65 2.29 2.64 HALC 1.40 1.54 1.63 1.65 2.15 1.67 2.24 2.69 VDGD (ours) 1.49 (+18%) 1.62 (+20%) 1.75 (+23%) 1.72 (+23%) 2.39 (+19%) 1.71 (+3%) 2.47 (+19%) 2.91 (+22%) MathVista Vanilla-greedy 1.56 1.65 2.00 1.92 2.56 2.15 2.45 2.97 Vanilla-sampling 1.54 1.68 1.91 1.94 2.52 2.19 2.46 2.95 VCD 1.67 1.68 2.07 2.11 2.59 2.53 2.81 3.19 OPERA 1.64 1.83 2.04 2.04 2.62 2.30 2.53 3.22 Woodpecker 1.72 2.10 2.19 2.13 2.43 2.43 2.59 3.13 LRV 1.68 1.68 2.17 1.99 2.61 2.20 2.62 3.30 LURE 1.59 2.03 2.21 2.07 2.37 2.45 2.49 3.13 PAI 1.68 1.85 1.97 2.08 2.63 2.20 2.60 3.39 HALC 1.71 2.05 2.17 1.97 2.54 2.24 2.77 3.38 VDGD (ours) 1.84 (+18%) 2.19 (+33%) 2.44 (+22%) 2.24 (+17%) 2.88 (+13%) 2.62 (+22%) 2.93 (+20%) 3.59 (+21%) MMBench Vanilla-greedy 2.67 2.89 3.11 2.98 3.58 3.19 3.71 3.93 Vanilla-sampling 2.69 2.92 3.12 3.02 3.59 3.22 3.68 3.95 VCD 2.80 3.03 3.09 3.22 3.93 3.28 3.81 4.20 OPERA 2.81 2.94 3.17 3.10 3.95 3.22 3.93 4.05 Woodpecker 2.83 2.96 3.14 3.07 3.84 3.33 3.83 4.19 LRV 2.77 2.99 3.13 3.23 3.84 3.25 3.77 4.05 LURE 2.72 2.97 3.13 3.26 3.72 3.21 3.82 4.16 PAI 2.75 3.00 3.04 3.16 3.94 3.35 3.97 4.05 HALC 2.73 2.93 3.02 3.18 3.81 3.27 4.04 4.08 VDGD (ours) 2.93 (+10%) 3.12 (+8%) 3.26 (+5%) 3.32 (+11%) 4.08 (+14%) 3.42 (+7%) 4.15 (+12%) 4.31 (+10%) MME Vanilla-greedy 3.32 3.54 3.65 3.49 3.75 3.57 3.64 3.86 Vanilla-sampling 3.34 3.53 3.62 3.48 3.77 3.58 3.66 3.85 VCD 3.46 3.61 3.77 3.59 3.96 3.67 3.95 4.21 OPERA 3.42 3.59 3.82 3.54 3.93 3.60 3.88 4.09 Woodpecker 3.37 3.55 3.76 3.63 3.82 3.59 3.92 4.15 LRV 3.43 3.63 3.78 3.52 3.95 3.61 3.96 3.99 LURE 3.42 3.62 3.87 3.67 3.93 3.72 3.77 4.10 PAI 3.38 3.56 3.81 3.55 3.89 3.69 3.82 4.04 HALC 3.47 3.58 3.83 3.71 4.01 3.59 3.80 4.14 VDGD (ours) 3.59 (+8%) 3.70 (+5%) 3.99 (+9%) 3.82 (+9%) 4.12 (+10%) 3.79 (+6%) 4.09 (+12%) 4.34 (+12%) Oven Vanilla-greedy 2.44 2.66 3.13 2.86 3.35 3.35 3.48 3.64 Vanilla-sampling 2.45 2.65 3.10 2.84 3.33 3.37 3.45 3.61 VCD 2.20 2.44 2.96 2.67 3.10 3.21 3.39 3.59 OPERA 2.50 2.49 3.20 2.92 3.44 3.42 3.53 3.68 Woodpecker 2.46 2.46 3.21 2.90 3.45 3.45 3.52 3.70 LRV 2.51 2.52 3.18 2.88 3.47 3.40 3.57 3.74 LURE 2.49 2.51 3.18 2.91 3.42 3.46 3.55 3.72 PAI 2.54 2.60 3.26 2.95 3.48 3.48 3.60 3.78 HALC 2.52 2.57 3.24 2.94 3.50 3.45 3.61 3.80 VDGD (ours) 2.78 (+14%) 2.84 (+7%) 3.49 (+12%) 3.15 (+10%) 3.68 (+10%) 3.59 (+7%) 3.75 (+8%) 3.92 (+7%) LLaVA-Bench Vanilla-greedy 3.01 3.12 3.23 3.13 4.12 3.68 4.15 4.37 Vanilla-sampling 2.98 3.10 3.27 3.17 4.08 3.71 4.17 4.33 VCD 2.93 3.16 3.18 3.15 3.89 3.57 4.13 4.28 OPERA 3.08 3.18 3.30 3.22 4.16 3.75 4.22 4.40 Woodpecker 3.10 3.22 3.36 3.23 4.21 3.73 4.21 4.39 LRV 3.16 3.19 3.30 3.22 4.27 3.72 4.20 4.49 LURE 3.18 3.26 3.35 3.24 4.25 3.64 4.27 4.51 PAI 3.20 3.28 3.38 3.25 4.30 3.74 4.32 4.54 HALC 3.25 3.27 3.37 3.22 4.24 3.76 4.35 4.53 VDGD (ours) 3.36 (+12%) 3.47 (+11%) 3.52 (+9%) 3.38 (+8%) 4.45 (+8%) 3.96 (+8%) 4.50 (+8%) 4.78 (+9%) MMVET Vanilla-greedy 2.39 2.52 2.72 2.87 3.39 3.18 3.47 3.64 Vanilla-sampling 2.34 2.50 2.70 2.84 3.32 3.35 3.42 3.62 VCD 2.29 2.61 2.67 2.92 3.29 3.42 3.39 3.58 OPERA 2.48 2.67 2.77 3.01 3.37 3.39 3.52 3.69 Woodpecker 2.47 2.64 2.82 3.06 3.40 3.44 3.50 3.72 LRV 2.58 2.56 2.79 2.92 3.38 3.42 3.51 3.75 LURE 2.54 2.63 2.81 3.12 3.41 3.46 3.56 3.71 PAI 2.62 2.60 2.83 3.18 3.37 3.49 3.59 3.78 HALC 2.66 2.70 2.88 3.15 3.44 3.47 3.54 3.76 VDGD (ours) 2.85 (+19%) 3.01 (+19%) 3.23 (+19%) 3.45 (+20%) 3.78 (+12%) 3.57 (+12%) 3.74 (+8%) 3.99 (+10%) VaLLu Vanilla-greedy 1.95 2.03 2.63 2.20 2.66 2.82 3.23 3.45 Vanilla-sampling 1.86 2.01 1.64 2.18 2.70 2.83 3.21 3.40 VCD 1.47 1.55 1.80 1.80 2.32 2.63 3.19 3.42 OPERA 2.04 2.05 2.62 2.26 2.65 2.85 3.22 3.47 Woodpecker 2.01 2.05 2.67 2.23 2.60 2.91 3.28 3.52 LRV 1.98 2.10 2.65 2.19 2.59 2.88 3.32 3.51 LURE 2.03 2.03 2.64 2.24 2.64 2.78 3.28 3.48 PAI 2.04 2.23 2.78 2.34 2.76 2.86 3.31 3.54 HALC 2.02 2.22 2.74 2.32 2.72 2.89 3.29 3.52 VDGD (ours) 2.16 (+11%) 2.64 (+30%) 3.16 (+20%) 2.72 (+24%) 3.45 (+30%) 3.01 (+7%) 3.48 (+8%) 3.67 (+6%)